OpenStudy (anonymous):

(-42x)(+2x)=84x .....is this wrong (-42x)(+2x)=84x .....is this wrong @Mathematics

OpenStudy (anonymous):

yes

OpenStudy (lukecrayonz):

Si.

OpenStudy (saifoo.khan):

Negative sign is missing. :O and x * x = x^2

OpenStudy (anonymous):

you will get -40x.

OpenStudy (anonymous):

That is wrong

OpenStudy (lukecrayonz):

No, remove the x on the second parenthesis, it will be right.

OpenStudy (lukecrayonz):

Right now you have -82x^2

OpenStudy (anonymous):

so it will be -84x^2

OpenStudy (lukecrayonz):

Sorry, -84x^2.

OpenStudy (anonymous):

-42*2=-84

OpenStudy (lukecrayonz):

Do (-42x)(2)=-84x :)

OpenStudy (anonymous):

thnx